OpenStudy (anonymous):

Factor. 56y^2 − 55y + 6

OpenStudy (abmon98):

(y+6/7)(y+1/8)

OpenStudy (abmon98):

solve quadratically

OpenStudy (anonymous):

thanks

OpenStudy (abmon98):

your welcome :)

OpenStudy (anonymous):

\[ 56 y^2 - 55 y + 6 = (-6 + 7 y) (-1 + 8 y) \]

OpenStudy (anonymous):

\[ \left(y+\frac{6}{7}\right) \left(y+\frac{1}{8}\right)=y^2+\frac{55 y}{56}+\frac{3}{28} \] @Abmon98, your answer is not right.

OpenStudy (abmon98):

its right but i just forgot to get rid of the denominators by multiplying it out.

OpenStudy (anonymous):

As written, it is not right.

OpenStudy (anonymous):

You do not have the polynomial =0 you only have a polynomial to factor